### v1

# Learning a Distributed Hierarchical Locomotion Controller for Embodied Cooperation

Abstract
In this work, we propose a distributed hierarchical locomotion control strategy for whole-body cooperation and demonstrate the potential for migration into large numbers of agents. Our method utilizes a hierarchical structure to break down complex tasks into smaller, manageable sub-tasks. By incorporating spatiotemporal continuity features, we establish the sequential logic necessary for causal inference and cooperative behaviour in sequential tasks, thereby facilitating efficient and coordinated control strategies. Through training within this framework, we demonstrate enhanced adaptability and cooperation, leading to superior performance in task completion compared to the original methods. Moreover, we construct a set of environments as the benchmark for embodied cooperation. A visual depiction of highlights of the learned cooperative behaviours can be found on this website [^1] Videos of our results are available at our project webpage: https://d-hrl.github.io.
Keywords: cooperation, locomotion, hierarchical reinforcement learning
1 Introduction
Cooperatively accomplishing embodied tasks by multiple robots has consistently been a highly challenging area of research. Recent studies mainly focus on embodied manipulation cooperation among robotic arms or formation control over the upper level within a group of mobile robots [1, 2]. Nevertheless, multi-agent cooperation via whole-body and end-to-end locomotion control is rarely studied.
Some previous works showcase the manipulation via locomotion [3] but are only tested on two agent systems, and the scalability of this method is still agnostic for migration to any number of agent populations. In this work, we aim to realize more complex embodied multi-agent cooperation by learning a distributed hierarchical locomotion control system, decomposing the complex and coupled behaviours while maintaining the potential for unlimited expansion on the swarm. As the foundation for implementation and validation, we construct three scenarios in IsaacSim/Gym [4] as benchmarks for embodied cooperation study.
(a) Cooperative Transport
(b) Corridor Crossing
(c) Ravine Bridging
Figure 1: Embodied Cooperation Environments.
Concurrently, training a robot for a specific function can be effectively achieved through reinforcement learning (RL), like learning movement patterns [5], interactive behaviours [6], as well as logical inference in games [7]. Although RL provides a recognized powerful exploration capability and tremendous progress has been made in sampling efficiency [4], finding and mastering a sequence of sophisticated tasks through searching remains a challenging problem. Hierarchical reinforcement learning (HRL) alleviates this to a certain extent, aiming to understand the logical relationships among ”control, action, behaviour, dynamic outcomes, and feedback” in a segmented manner. An increasing number of studies are employing hierarchical approaches to decompose complex tasks [8, 9, 10, 11, 12, 13, 14, 15]. Inspired by these, we propose a three-level HRL to naturally decompose perception feature extraction, cooperative behaviours, and fundamental actions.
In the previous work [3], embodied cooperation is completed via locomotion control by two quadruped robots. However, spatiotemporal features between a sequence of events are not considered in this work, which causes failure in cooperation of more than two agents. Actually, understanding interactive and cooperative behaviours requires sequential causal reasoning considering world dynamics. To enhance this, we deploy sequential behaviour learning with spatiotemporal memory capabilities in the middle layer of the hierarchical network by constructing a Recurrent Neural Network (RNN). Furthermore, the previous method can hardly be applied to a flock of agents [3], while the scalability of the swarm is crucial in multi-agent reinforcement learning (MARL) field [16, 17]. Herein, we construct our method facing the decentralized and distributed scenario and attempt to apply it to an expandable MAS.
The key contributions of this work can be described as follows:
  • A distributed HRL locomotion control framework is constructed for scale-expandable embodied cooperation. This architecture can be applied to both homogeneous and heterogeneous groups.
  • We build a set of scenarios for whole-body embodied cooperation in IsaacSim shown in Fig. 1, as a benchmark for embodied cooperation.
  • We showcase the remarkable emergence of cooperative behaviours among agents and demonstrate the advantages compared with previous approaches.
Figure 2: Distributed HRL pipeline: demonstrate the information flow within the distributed hierarchical reinforcement learning, using scenario Cooperative Transport as an example. Here we maintain a centralized training but distributed control hierarchical reinforcement learning framework, where we decompose the complex behaviours into three levels: Upper Layer (UL), Middle Layer (ML), and Lower Layer (LL). UL module processes the external information $e_{t}$ including environmental perception and relative position to colleagues, extracting features and sending to ML; ML module is a recurrent neural network layer that maintains a recurrent state $h$ considering temporal and spatial correlation, and outputs a locomotion command into LL module; LL module is a pre-trained locomotion control layer that generates action $a_{t}$ and applies it to agent according to the proprioceptive observation $p_{t}$, where we have two modes: position and velocity. The goal of Cooperative Transport is to move a cylinder object to the red target zone collaboratively by a group of Ant robots.
2 Related Works
2.1 Embodied Cooperation
Research on cooperation can be divided according to with or without body interaction. Some collaborative tasks are based on particle games, focusing solely on coordination and alignment at position/velocity levels [18, 19]. In contrast, studies related to embodied cooperation lie on the foundations of the robotic platform that owns the real body, paying more attention to interaction, collision, and whole-body control. Optimization methods and computational intelligence methods can solve embodied cooperation problems. The model predicted control (MPC) based methods are widely used in formation control on quadruped robots [20, 21, 22, 23] and aerial robots [24, 25], and nature-inspired algorithms can obtain good performance in a self-organized swarm cooperative task [2]. These approaches provide highly accurate control but lack flexibility and generalization. Learning-based methods can compensate for this [26, 3]. Moreover, the large language model (LLM) can provide robots with powerful cognitive reasoning capabilities for collaboration, bringing a new way to solve the problem [27, 28, 29, 30].
2.2 HRL in locomotion Control
HRL is well suited for complex and spatiotemporal tasks [31]. Its effectiveness has been proven and widely acknowledged [32]. Multiple steps can decouple an intricate task to reduce the training difficulty, corresponding to various layers of policies [33]. Decisions between layers can be made by deciding whether the previous action has been completed [34]. Essentially, hierarchical learning provides artificial direction and reduces the need for model to have a thorough understanding of the scenario and the tasks, preventing RL from wandering in a sparse search space.
Furthermore, HRL is inherently well-suited for tasks that involve a progressive logical relationship [35, 36]. A typical work fuses simplified environmental observation and proprioceptive state then generates lower-level control commands for actuators[37]. Some recent attractive benchmarks like HumanoidBench [38], also divide tasks into planning and reaching control and learning hierarchically.
3 Problem Statement
We first formulate the problem of learning embodied cooperative behaviour in a swarm of agents, then try to solve this problem with MARL. Since the number of agents is fixed within an episode, the synchronous framework and policy gradient descent algorithms are selected. To realize real-time execution, we further select the Centralized Training and Decentralized Execution (CTDE) framework to support our implementation. It is crucial to mention that agents in the group might be homogeneous or heterogeneous, and agents in the same species share the same policy.
The embodied cooperation problem can be formulated as a Markov Decision Process (MDP) mathematically expressed as $\left\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R}\right\rangle$ including $N$ agents in $M$ heterogeneous species. $\mathcal{S},\mathcal{A}$ are state and action tuples, respectively, where $\mathcal{S}=(S_{1},S_{2},\cdots,S_{N})$, $\mathcal{A}=(A_{1},A_{2},\cdots,A_{N})$. $\mathcal{T}$ is the state transition function, and $\mathcal{T}\left(s^{\prime}|s,a\right)$ represents the transition probability distribution of the system when transiting into the next states $s^{\prime}=(s^{\prime}_{1},s^{\prime}_{2},\cdots,s^{\prime}_{N})\in\mathcal{S}$ from states $s=(s_{1},s_{2},\cdots,s_{N})\in\mathcal{S}$ after taking a set of actions $a=(a_{1},a_{2},\cdots,a_{N})\in\mathcal{A}$. Besides, $\mathcal{R}=(R_{1},\cdots,R_{M})$ denotes the tuple of reward functions in which different types of agents correspond to distinct reward functions. $r_{i}={R_{i}}\left(s_{i},a_{i}\right)$ is the reward given to agent $i$ after taking actions $a_{i}$ under states $s$.
For the homogenous multiple agents group, the reward function $R$ is determined and shared for each agent. The aim is to find the best shared policy $\pi^{*}$, to maximize the expected average value of cumulative discount reward:
[…]
where $s_{t}^{i}\in S_{i}$ and $a_{t}^{i}\in A_{i}$ are the state and action of the $i$-th agent. There are some subtle differences in the representation of the heterogeneous population. The aim is to find the best policy tuple $\Omega^{*}=(\pi_{1},\cdots,\pi_{M})$ to maximize the joint multiplication of costs $J$:
$\Omega^{*}=\mathrm{arg}\underset{\Omega}{\mathrm{max}}\prod_{m=1}^{M}J_{m},\;\mathrm{where}\;J_{m}=\frac{1}{K}\sum_{i=1}^{K}\sum_{t=0}^{\infty}\gamma^{t}\cdot\mathbb{E}\left[R_{m}(s_{t}^{i},a_{t}^{i})|\pi_{m}\right]$ (2)
where $J_{m}$ denotes the cost function of type $m$ species, and $K$ is the agent number of this species. The final optimization cost for the heterogeneous population is a form of co-multiplication because task assignment exists among the entire team. Each subtask must be completed otherwise the overall task will fail. Even if some subtasks obtain high rewards, any failure in a subtask will prevent the achievement of the overall goal.
4 Distributed Hierarchical Reinforcement Learning
As aforementioned, locomotion control for embodied cooperation is an extremely challenging task. There are three main difficulties: (i) How to decompose and decouple the complicated tasks into a step-by-step mode? (ii) How to establish the temporal and spatial correlation of behaviours? (iii) Whether the trained models can expand to a large scale of multiagent collaboration? In this part, we will address each of these three questions in detail.
4.1 Task Decomposition and Hierarchical Learning
Network Structure: Traditionally, a hierarchical neural network structure divides tasks into control and decision layers. Unlike this, in order to address collaborative tasks, we propose a deeper hierarchical structure consisting of three layers. We designate perception-related tasks as the upper layer of the three-layer network that extracts exteroceptive observation including other agents’ information into a feature vector; then, we define the middle layer of the network as the planning and decision layer, which provides spatiotemporally correlated commands to the lower-level controller; finally, the lower-layer network focuses on the execution of commands, which can be pre-trained in various modes. The overall information flow in the proposed hierarchical network is illustrated in Fig. 2.
Hierarchical Perception Processing: Each layer of the network is responsible for processing distinct types of raw data, which relates to how we acquire and classify perceptual information. Proprioceptive perceptual states including local position, velocity, and joint torque are closely associated with personal locomotion control. Therefore, the proprioceptive state $p$ should only processed by the lower-layer policy. Furthermore, the exteroceptive state $e$ formed by concatenating environmental perception which provides terrain information in the form of a height measurement array [39] with relative observations of other agents [5], as depicted in Fig.3.
Figure 3: Exteroceptive state configuration in distributed HRL.
4.2 Spatiotemporal Memory Recurrence
Capturing and leveraging spatiotemporal features is significant for long-term decision-making tasks [40]. For tasks requiring cooperation, spatial awareness is as crucial as understanding temporal dynamics. We propose spatiotemporal memory recurrence that uses RNNs to enhance the agent’s ability to make informed decisions over extended periods and complex environments.
The significance of learning spatiotemporal features through RNNs in distributed HRL lies in the inherent capability of these networks to maintain and utilize historical information. This capability is particularly important in dynamic environments, especially in nonstationary multiagent environments. RNNs, on the other hand, are designed to handle sequential data, making them ideal for tasks that require an understanding of the temporal progression of states and actions. Thus, we maintain a hidden recurrent state $h$ to pass the feature of scenarios continuously, as shown in Fig. 2.
4.3 Distributed Scalable HRL
In contrast to previous studies that employed HRL or end-to-end approaches for robotic cooperative manipulation or locomotion, our distributed HRL method considers the capability of scaling the population. To achieve this, we need to strictly define the format of input and output information for each agent, particularly the observation. Herein, we propose to use partial observation of agents related to their colleagues, according to their visibility [41]. Each agent can only obtain the nearest neighbour’s local position. As illustrated in Fig.2, the dimension of exteroceptive state $e$ is limited. This actually changes the difficulty of collaboration learning. Since partially observed agents can only see nearby peers and environmental information, they need to estimate the states of other invisible agents from the dynamics of the environment and predict future changes in the environment.
4.4 Training Curriculum
Acquiring collaborative behaviours requires the agents’ random search to span a sufficiently large space, which in complex tasks can often lead to an inability to learn usable knowledge due to the vastness of the search space. In this case, we design some appropriate training curricula that reasonably use sparse rewards and dense rewards in the early stages to intentionally guide the agent to approach success. This curriculum is independent of HRL pertaining, such as encouraging agents to move to the target first at the early stage. The curriculum designs are described in detail in the appendix.
5 Experiments
The experiments are conducted on a workstation with Intel Xeon W-2150b and RTX2080Ti. Independent Proximal Policy Optimization (IPPO) in a multi-agent setting is employed as the algorithm. Adam optimizer is used with a learning rate of 0.0005. PPO clipping is 0.2, the discount factor is 0.995. We could complete the training within several hours thanks to IsaacSim GPU parallel sampling via multiple environments [4].
5.1 Environments Construction
Three scenarios are established that represent different task types: Cooperative Transport, Corridor Crossing, and Ravine Bridging.
Cooperative Transport: Transporting an object collaboratively is a classical task in the multiagent system field. Herein, we build an embodied cooperative transport task via locomotion control. A group of Ant attempt to move a cylinder object to the red target area, shown in Fig.1(a). We consider fully decentralised and distributed cooperation in a homogeneous and self-organized flock. Thus, the scale of flock can have theoretically unlimited expansion.
Corridor Crossing: A group of agents navigate through a narrow corridor sequentially, aiming to complete the passage as quickly as possible. Only one person can pass through the slit at any one time. Agents in this scenario should adjust their position and speed reasonably according to the surrounding observations and pass through a narrow passage without collision. Cooperation in formation is required, as shown in Fig. 1(b). The number of agents can also be expanded in this task.
Ravine Bridging: As illustrated in Fig. 1(c), two agents push a movable bridge, enabling another group of agents to cross the chasm smoothly via the bridge. There are two categories of agents in this scenario, which is different from the other two environments. Agents in Ravine Bridging are heterogeneous, which means we should maintain more than one policy for different missions.
5.2 Results of Distributed HRL
5.2.1 Training Basic Locomotion Control
In the training of lower-level motion control, two control paradigms are available: velocity control and position control. We first train the two control paradigms in the single-agent environment to obtain accurate control of the fundamental actions. We give a reciprocal $L_{2}$ distance between the current position and the target position, or the dot product of the agent velocity and target velocity, as the main reward in each training curriculum. Fig 4(a) shows the training results of two paradigms. We find that the velocity control offers greater flexibility than position control. In our empirical tests, however, position control presents higher accuracy. Moreover, velocity control makes it difficult to maintain a stationary position at a specific point. Similar to any other hierarchical learning study, the lower layer is frozen once its training is done well.
(a) Basic Locomotion
(b) Cooperative Transport
(c) Corrider Crossing
(d) Ravine Bridging
Figure 4: Distributed HRL results for all scenarios.
5.2.2 Training the Interactive Collaboration
After the lower layer is elaborately trained, we freeze this network and fix the basic locomotion control performance. Then, the upper and middle layers are trained on different scenes via distinct reward mechanisms. The training results are shown in Fig. 4(b,c,d).
[…]
Ravine Bridging: The group in this task is a heterogeneous team. Although all agents are Ant, there are two sets of policies: one for agents on the plane and one for agents in the ravine. Agents on the plane will arrive at the edge of ravine and stop, waiting for the agents in the ravine to transport the bridge to the front of them. We get approximately half success over one hundred trails. The main reason is that Ants in the ravine might get stuck or fall.
Table 1: Success Rate Comparison: We present success rates with standard error on the three tasks, comparing our method to ablations. Each percentage results from 50 trials—10 attempts per three independently trained models.
Success rate HRL No Hierarchy No Spatiotemporal Memory
Cooperative Transport $\textbf{74}\pm 7\%$ $0\pm 0\%$ $5\pm 1\%$
Corridor Crossing $\textbf{93}\pm 3\%$ $0\pm 0\%$ $2\pm 1\%$
Ravine Bridging $\textbf{48}\pm 5\%$ $0\pm 0\%$ $0\pm 0\%$
Figure 5: Illustration of hierarchical learning concepts applied in different scenarios compared with no hierarchy case, noisy hidden states case and the baseline [3] which does not have the middle layer with recurrent neural networks. All results are conducted on four agents scenario.
5.3 Ablation
5.3.1 The role of Hierarchical Learning
In this experiment, we replace the hierarchical networks with a single-layer MLP, shown in Fig. 5. We find the reward converges to an extremely low level quickly. Agents can only learn and get the basic reward of locomotion. These results show that hierarchy is crucial to success in more complex tasks that require coordination. Without hierarchy, agents cannot complete the mission (in Table. 1). The reason for failure is obvious. As with other HRL objectives, a single network facing high-dimensional observations and complex multi-item reward functions finds it difficult to distinguish features between channels and learn behaviour guided by a single main reward.
For multi-agent training, the situation is even more unique. Due to the CTDE training paradigm we default to, differences in the positions of multiple bodies lead to an inability to uniformly learn certain distance rewards. For example, in the task of cooperative transport, agents have different positions compared to objects and targets, so a unified strategy cannot be directly formed by lower-level networks. It can be said that hierarchical learning is indispensable for such group tasks.
5.3.2 The effectiveness of Spatiotemporal Memory
To validate the effectiveness of spatiotemporal features, we re-product the method in previous work [3] as the baseline, which does not consider the event correlation. As shown in Fig.5, our method obtains a much higher reward for all three tasks. The approach proposed in  [3] which does not have RNNs, converges to a slightly higher reward than in the case of no hierarchy but is still far from completing the mission. Furthermore, we attempt to replace the RNN hidden states with random noise inputs to explain how event correlation positively and effectively affects the entire performance. The success rate of all these attempts is close to zero. We find that agents cannot learn a set of systematic behaviours in specific tasks, and there are occasional successes when the object position is really close to the target. Actually, if the agents only react to the current observation, they could never form strategic behaviours. In other words, he would never think for himself in the next second without spatiotemporal memory. This is even more deadly in collaborative tasks, as we can obviously find that the conciliatory and sacrificial behaviour emerging in crossing corridor tasks is a strategy of thinking for long-term consequences.
5.4 Analysis
Scalability Analysis: We validate scalability in the task cooperative transport as illustrated in Fig. 6(a). The final reward increases as the number of agents increases, which also aligns with our expectations for this task. More agents participating should indeed bring better transportation results. Our results reveal that the distributed HRL has great potential to migrate a trained policy to a large number of agents. Only such algorithms can achieve the true sense of ”biocoenosis” collaborative tasks.
(a) Training results for different Ant populations
(b) Distributed HRL
(c) No hierarchy
(d) No memory
Figure 6: (a) shows the training results of different Ant populations used in Cooperative Transport. (b)(c)(d) shows the success and failure distribution of task Cooperative Transport.
Typical Result Analysis: We analyse the typical result of task cooperative transport. To test the robustness of our method, we conduct more than one hundred trials with randomized reset positions and count the success rate then plot the scatter diagram, as illustrated in Fig. 6(b). Each point represents an initialized position of an agent, and the object is located at (0,0) point. Red points denote agents generated at these positions fail to complete the task, while each blue point denotes a success. We can observe that when the initial position is closer to the object, the success rate tends to be higher, and there is only one failure case within a 5-meter range of the object; conversely, the success rate decreases as the distance from the object increases. This is because the probability of accidents is greatly reduced by short-distance tasks. It is worth mentioning that most of the failure cases result from the agent tipping over or pushing the object obviously to the side with uneven force during the pushing process. Fig.6(c)(d) showcase the evaluation results under no-hierarchy and no-spatiotemporal memory. Only few successes can be found in the evaluation of no-spatiotemporal memory, and all cases fail in training without hierarchy.
6 Conclusion
In this work, we propose a distributed HRL architecture for embodied cooperation, and the results showcase the effectiveness of our method in embodied and whole-body control collaborative tasks. Besides, our approach has significant potential in scaling up the flock size. In the future, more complex cooperative tasks can be conducted based on our method.
Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
Appendix A Appendix
A.1 Environment Description
A.1.1 Cooperative Transport
The agents should cooperate to push a cylinder to a target location. The local positions between the agents, cylinder, and target are initialized randomly. The cylinder is set in the middle of the map, and agents are initialized in the four quadrants of the cylinder respectively.
Observation The agents receive two sets of observations: (i) a set of ”proprioceptive” states which would be given to the lower layer controller. This part of observation is exactly the same in all agents in different tasks because all tasks use the same pre-trained lower layer policy, including their self-observed information, such as its local position, local velocity, angular velocity, orientation, body upright angle, joint force and its limitations, etc. (ii) a set of “exteroceptive” features containing task-relevant information which would be given to upper layer controller. In this task, the observation includes the locations of the nearest agent, the current position of the cylinder, and the destination.
Rewards We set three types of rewards: (i) Cylinder distance reward: a term which is the reciprocal L2 distance of the agent to the target. This is the main reward of this task. It will increase when the cylinder becomes closer to the target. The increasing trend is the same as the inverse proportional function. (ii) Agent cylinder distance reward: this term is to encourage agents firstly come closer to the cylinder so that they can learn how to push the object. (iii) Agent distance reward: a term used to guide the agent to walk towards the target, thereby obtaining a large cylinder distance reward.
A.1.2 Corridor Crossing
The agents are expected to reach a designated line through a narrow slit. Two walls block most of the way forward, leaving only a narrow corridor that only allows one agent to pass at a time. Yet all agents are at the same distance to the entrance of the slit, which means that agents would get stuck at the entrance of the slit together if nobody is slow or stop to wait for others to pass first. We expected agents can acquire a strategy so that all the agents could go through the corridor as fast as possible.
Observation The lower layer controllers’ observation is the same as the last task. In this task, agents’ upper layer controllers’ observation includes the locations of itself and the nearest neighbour, the position, translation, and orientation of the walls.
Rewards Two rewards are set: (i) Destination reward: if an agent has passed the finish line, all agents will receive a positive reward. It stimulates the strategy allowing as many agents as possible to pass through the corridor to the finish line. (ii) Distance reward to enter: a term to encourage agents first to come close to the corridor.
A.1.3 Ravine Bridging
The agents are expected to reach a designated line where they must cross over a ravine first. There is a ravine blocking the way of the agents. Agents cannot jump over or climb out of the ravine. Agents need colleagues’ help to come over the ravine. At this point, we place a stone of the right size and depth in the ravine to serve as a ”bridge” to help the agents cross it. We pre-set two agents on either side of the stone to push the stone to help their colleagues cross the ravine and reach the destination. So this is a complex task that requires teamwork among heterogeneous agents.
Observation The observation of the lower layer policy is similar to the previous tasks. Agents’ upper layer controllers’ observation includes the locations of itself and the nearest agents, the position of the stone, and the position and depth of the ravine.
Rewards Two rewards are set: (i) Destination reward: if any agent has passed the finish line, all agents will receive a positive reward. It stimulates the strategy allowing as many agents as possible to go through the ravine and get to the finish line. (ii) Bridging reward: a term to encourage agents in the ravine to carry the stone to the agent above which is the closest to the ravine.
A.2 Curriculum
We utilize sparse rewards to direct agents towards proximity with the object, subsequently acquiring the target’s location in the Cooperative Transport task. In the Corridor Crossing task, we train individual agents to navigate to the finish line by passing through the slit and then refine the pre-trained policy for application in a multi-agent environment. The learning sequence for the Ravine Bridging task is more intricate which means we have to give more guidance; we instruct agents to wait at the ravine’s edge at the opportune moment to prevent falling, while other agents within the ravine initially learn to convey ”bridge” to an appropriate location.
A.3 Scalability
The distributed HRL framework allows the collaborative tasks to be explicitly represented as a ”group”, thereby enabling tasks to be extended to large numbers of agents team. We have tested the framework with different populations of Ant in various tasks (There are only two Ants for pushing bridge in the task Ravine Bridging), and the success rates of each task with different populations are shown in the Tab. 2. The results show the potential of our framework for handling tasks with more agents.
Table 2: Each percentage results from 50 trials—10 attempts per three independently trained models.
Success rate Two Ants Three Ants Four Ants Five Ants
Cooperative Transport $8\pm 2\%$ $42\pm 3\%$ $74\pm 7\%$ $82\pm 5\%$
Corridor Crossing $98\pm 2\%$ $95\pm 2\%$ $93\pm 3\%$ $79\pm 2\%$
Ravine Bridging $48\pm 5\%$ $36\pm 8\%$ $30\pm 2\%$ $12\pm 5\%$
Table 3: Success Rate Comparison between average moving distances in task Ravine Bridging.
Distance from Ravine Two Ants Three Ants Four Ants Five Ants
2 m $49\%$ $45\%$ $38\%$ $35\%$
4 m $38\%$ $33\%$ $20\%$ $11\%$
6 m $30\%$ $19\%$ $5\%$ $3\%$
A.4 Typical Result
A.4.1 Corridor Crossing
We record the success rates of each Ant at its initial position separately, as illustrated in Fig. A.4.1. We have a higher success rate when the initial position is closer to the center, and a higher failure rate when the initial position is closer to the edge. This is also explainable, as the Ant at the centre is closer to the slit and does not need to adjust velocity.
A.4.2 Ravine Bridging
The bridging and coordination in the ravine is the key to completing the task. We set up three difficulty levels - easy, medium, and hard - based on the distance from the ravine’s edge. The easy level corresponds to a distance of 2 m, the medium level to 4 m, and the hard level to 6 cm. The Tab.  3 shows the success rates of the three difficulty levels with different numbers of Ants. We can observe that it becomes more difficult for any number of Ants to complete the task as the distance increases. Because cooperation on the timing of stopping and bridging becomes more important when moving length increases.